Mirage Mirror's predetermined overhead rate for manufacturing overhead is $20 per direct labour-hour. The direct labour rate is $25 per hour. If the budgeted direct labour cost was $550,000, what was the budgeted manufacturing overhead?
Answer #1

Budgeted Number of hours = Labor cost / labor rate = 550,000 / 25 = 22,000 hours

Predetermined overhead rate = Budgeted manufacturing overhead / Budgeted direct labor-hours

$20 = Budgeted manufacturing overhead / 22,000

Budgeted manufacturing overhead = 22,000 * 20 = $440,000